Work Interruptions Survey found that out of 300 workers, 245 said they were interrupted three or more times an hour by phone messages, faxes, etc. Find the 90% confidence interval of the population proportion of workers who are interrupted three or more times an hour.


1
Expert's answer
2022-05-10T18:50:23-0400

The sample proportion is computed as follows, based on the sample size "N = 300" and the number of favorable cases "X = 245:"


"\hat{p}=\dfrac{X}{N}=\dfrac{245}{300}=\dfrac{49}{60}"

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}})"

"=(\dfrac{49}{60}-1.6449\sqrt{\dfrac{\dfrac{49}{60}(1-\dfrac{49}{60})}{300}},"

"\dfrac{49}{60}+1.6449\sqrt{\dfrac{\dfrac{49}{60}(1-\dfrac{49}{60})}{300}})"

"=(0.7799, 0.8534)"

Therefore, based on the data provided, the 90% confidence interval for the population proportion is "0.7799< p < 0.8534," which indicates that we are 90% confident that the true population proportion "p" is contained by the interval "(0.7799, 0.8534)."